# Measurement of the scintillation resolution in liquid xenon and its impact for future segmented calorimeters

### v5

@@ I INTRODUCTION
All of the above crystals have either been used in, or actively pursued for, high energy and nuclear physics experiments. NaI was used in Crystal Ball, CsI in CLEO and BGO in L3. LYSO is one of the best candidates for future collider applications, due to its excellent properties (high light yield, fast decay time, high density) and radiation hardness [5]. LYSO is also the preferred scintillation material in modern Positron Emission Tomography (PET) scanners [6, 7, 8]. PET scanners also benefit from a good energy resolution, because energy resolution allows one to reject a larger Compton scatter fraction, i.e., coincidences where one or both photons interact in the body of the patient and lose part of their energy. Those coincidences add noise to the image, since the direction of the gamma changes after the scattering.
@@
The common element in many nuclear and particle calorimeters and in modern PET scanners is a Segmented Scintillating Block, SSB (Fig. 1). An SSB is an array of contiguous cells filled with scintillating material. The light propagates inside each cell by reflection, until it reaches the readout sensor, normally a SiPM in modern devices, although PMTs and APDs were used in the past. The thickness of the cells depends on the application and varies between one and a few attenuation lengths. The SSB can be characterised by three parameters: energy resolution, which depends on the scintillator’s light yield, as well as the optical collection efficiency of the cell, the photodetection efficiency (PDE) of the SiPM, electronic noise, etc; time resolution, which is dominated by the scintillator’s decay time; and (transverse) spatial resolution, which can be tailored by choosing a suitable transverse size for the cell. As an example, in the 2 m long EXPLORER scanner [9, 10], the largest reported PET, the SSB material is LYSO, with cell area mm2 and thickness of 1.6 attenuation lengths, and the cells are read out by SiPMs. The number of individual cells exceeds . The energy resolution at 511 keV is 12% FWHM, the spatial resolution 3 mm and the time-of-flight (TOF) resolution is 500 ps.
@@
Liquid xenon (LXe) is also an excellent calorimeter [11]. Charged particles interact with LXe via both ionisation and excitation. The excitation of xenon atoms results in strongly bound excited molecules (excimers) which, on de-excitation, emit VUV photons with a wavelength of 175 nm [12]. In addition, the xenon atoms are ionised. An external electric field can be used to carry away the ionisation electrons, in which case two signals are available, one from scintillation, and one from ionisation. In the absence of such electric field, most ionisation electrons eventually recombine with positive xenon ions, and both channels (primary scintillation and recombination) contribute to the resulting scintillation yield of photons per MeV [13], which is a factor 1.5 larger than NaI and 1.7 larger than LYSO. Several measurements have found yield fractions of primary scintillation and recombination of roughly 0.26 and 0.74 for electrons of around 1 MeV of energy [14]. The primary scintillation is characterised by two decay constants ( ns, ns), while the decay time associated with recombination is ns [15]. Thus, LXe is, a priori, a brighter and (slightly) faster scintillator than LYSO. However, its attenuation length is almost three times longer than that of LYSO and the fraction of photons which interact via the photoelectric effect (relevant for PET application) a factor of two less.
@@
LXe has been used extensively in particle and astroparticle physics. In particular, the MEG experiment [16] deployed a large calorimeter —800 litres volume filled with LXe—, with a thickness of 17 attenuation lengths and read out by 846 UV-sensitive photomultiplier tubes of 2-inch diameter. Other important examples are experiments searching for rare events, in particular neutrinoless double beta decay [17] and direct Dark Matter searches [11, 18]. Future experiments such as nEXO [19] and DARWIN [20] are also based on LXe. All these detectors are Time Projection Chambers, able to read the scintillation and ionisation signals in LXe. Since the energy deposited by ionising particles is distributed between both channels, the anti-correlation of the signals leads to a good energy resolution. For example, the XENON1T experiment measures 0.8% at 2.46 MeV [18].
@@
The possibility of building a PET scanner based on LXe was proposed more than four decades ago [21] and explored by a number of subsequent works [22, 23, 24, 25]. The advantages of LXe compared to scintillating crystals are a large light yield, a fast decay time and the fact of being a liquid medium, which reduces the complexity of the system and increases its scalability. The disadvantage of having to add a cryogenic system to keep xenon in its liquid phase (the liquefaction temperature of xenon is -112∘C at 1 bar) is not problematic, since cryogenic systems reaching much lower temperatures are already in use in hospitals, for instance, to cool down magnets in magnetic resonance scanners. Two strategies were considered: a PET scanner based on scintillation only, and one that would combine both ionisation and scintillation. Given the very high rate of PET scans, the second strategy may be unpractical, due to the long time (microseconds) needed to drift the charge. The LXe micro-PET prototype obtained an energy resolution of 4.3% [25] combining scintillation and charge, while the XEMIS1 prototype measured [26] with charge only. On the other hand, several groups tried prototypes based on pure scintillation (measured by PMTs), obtaining an energy resolution of 6.4–13.6% [24, 25, 27]. None of these results were compelling enough to consider LXe as a viable alternative to standard inorganic crystals, in particular LYSO. Recently, the possibility of building a LXe-PET using modern SiPM technology (thus providing better energy resolution and time-of-flight than that possible in the past) has been considered by our group [28, 29, 30, 31, 32, 33, 34].
@@
In 2002, Doke studied the intrinsic resolution of LXe [35] and found that non-proportionality of the scintillation yield resulted in % in the region between 100 keV and 5 MeV (Fig. 2). His calculation is in good agreement with the modern Monte Carlo code describing the physics of LXe, NEST [36]. Specifically, at 511 keV, NEST predicts % .
Several experimental groups have reported energy resolution measurements in LXe for different gamma-rays, but none of these obtained the intrinsic resolution. For instance, a work from Yamashita et al. obtained an energy resolution of 12 at 122 keV and 7.1 at 662 keV [37], while Aprile et al. reported 5 at 122 keV [38]. The only attempt to go beyond simple energy resolution is that of Ni and collaborators, who estimated a combined contribution of intrinsic resolution and electron-ion recombination fluctuations, obtaining 6-8 [39] for gamma rays between 122 keV and 662 keV. Although this value is not directly comparable to the result obtained by Doke, this result is almost a factor 5 worse.
In this paper, we present a new measurement of the energy resolution in LXe, using scintillation only. A value of the intrinsic resolution combined with the fluctuation due to recombination is also shown. Our setup is conceptually identical to the SSB described above, and illustrated in Fig. 1, where the scintillating material is now LXe. The advantage of the SSB is its large light collection efficiency, as well as the availability of VUV-SiPMs with large PDE. This results in a large number of detected photoelectrons (pe), thus minimising Poisson fluctuations and leading to a measurement close to .
@@ II SETUP
The active region of the setup consists of two LXe SSBs housed in a metal-sealed aluminum cube (F100-1ALU) from Vacom. The xenon is continuously recirculated in gas phase by a double diaphragm compressor and flows through a model PS4 MT15 R2 hot getter from Sigma Technologies. This process removes nitrogen, oxygen and water from the xenon, all of which quench xenon scintillation light [40]. The internal cube is housed within a large vacuum vessel so that the structure is thermally isolated while cooled. Said cooling is performed using a Sumitomo CH-110 cold head coupled to the cube via custom-made copper thermal links. Heat dissipation from the thermal links is minimised by wrapping them in polyethylene terephthalate foil. The cube and cooling machinery is visible in Fig. 4. The gas passes through a heat exchanger when entering and exiting the inner cube, which precools the gas before liquefaction for operation at 1.26 bar.
@@ III MONTE CARLO SIMULATION
A detailed simulation of our set-up has been developed using a Monte Carlo software application based on Geant4 [43]. The physical and optical properties of the materials in use are simulated thoroughly; in particular, the reflectivity of PTFE has been modelled as a Lambertian reflectivity with a value of and the value of the refractive index as a function of the wavelength has been extracted from Ref. [44]. The other physical parameters relevant for optical photon propagation are the Rayleigh scattering length (36 cm [45]) and the transparency and refractive index of the quartz window placed in front of the SiPMs to protect them (90 and 1.6, respectively). The photodetection efficiency provided by the sensor datasheet is also included as a function of the wavelength of the optical photons. We turn off any intrinsic fluctuations of the LXe response to gamma interactions, so that the amount of scintillation photons produced is proportional to the deposited energy.
The simulation of a 22Na source placed in the calibration port provides a charge distribution with the photopeak centred in 5240 pe. Since the number of microcells in our SiPMs is 6162, it is likely that our measurement is affected by saturation. In order to confirm that, we ran a second simulation, where each microcell of the SiPMs is simulated and the recovery time of our sensors (158 ns at -110∘C) taken into account, to model the loss of photons due to saturation; we found that in this case the photopeak is centered in 3752 pe, which implies a significant charge loss. The same simulations, with and without saturation effects, were carried out for two other available calibration sources, with lower energy gammas, namely 133Ba (provides gamma lines at 30.8, 81 and 356 keV) and 57Co (provides gamma line at 122 keV). In Fig. 6 the number of fired microcells versus the number of photons that should be detected in the case of negligible saturation is shown for MC events at all energies. The saturation curve provided by Hamamatsu [46] is superimposed on top of this distribution, showing the excellent match with the simulation. The positions of the photopeaks of the three calibration sources lie close to the Hamamatsu model.
@@ IV MEASUREMENT
The data acquisition output is a list of channels with integrated charge and corresponding timestamps. Once the channels are divided event by event, a first filter is applied to retain only coincidences, i.e., events where at least one sensor from each detection plane is present. This condition is applied only to data taken with the 22Na source. Given the cell structure of the SSBs, most of the charge is read out by the sensor coupled to the xenon volume where the interaction occurred, therefore, we single out the sensor with larger detected charge for each plane. The same procedure was applied to the 133Ba and 57Co sources.
Once the energy spectrum was obtained for the three sources, the ASIC effect due to the integrator was corrected and the energy spectrum in pe was obtained (see Fig. 7).
@@
where and are the best values for the Gaussian parameters coming from the fit. This procedure is applied to all the sensors in the SSB and their results are combined using a weighted mean, where each value is weighted by the inverse of its squared fit error (i.e., more precise measurements have a larger influence on the final resolution), thus yielding a final resolution of . The total uncertainty is calculated as the combination of two contributions: the statistical uncertainty from the fits and the spread across sensor resolutions.
The energy resolution is slightly better than the value predicted by MC without SiPM saturation, which is (see Fig. 9). The MC error includes the uncertainties in the optical modelling, such as the reflectivity of the teflon and the reflection and refraction between LXe and the quartz window that protects the SiPMs (see Fig. 3–bottom).
The low value of the energy resolution compared to MC is due to SiPM saturation, which has the effect of compressing the data and brings on an artificial better energy resolution. Therefore, correcting the data with the saturation curve provided by Hamamatsu, an energy resolution of is obtained, showing compatibility with the MC result within . Fig. 10 shows the overlaid energy spectra of the experimental and MC data without saturation showing a good agreement. Fig. 11 presents the corrected energy resolutions for all the SiPMs in the reference SSB.
@@ V RESULTS
The energy resolution obtained in our work is shown in Fig. 12, together with others’ measurements. It can be seen that our results are 2-3 times better than previous works using only scintillation. The only slightly better measurement is the one from Reference [38], which only measured one peak.
Following Doke [35], the measured resolution, , can be seen as the quadratic sum of the intrinsic resolution, , the resolution due to Poisson variation in photon detection, and the resolution coming from the fluctuations of electron-ion recombination due to escape electrons, :
@@
Since our MC simulation does not include neither the intrinsic nor recombination fluctuations, the value we extract substracting the measured square resolution and the MC square resolution provides us with the combination of these two resolutions, which we call .
@@ VI DISCUSSION
The combined intrinsic resolution that our procedure yields at 511 keV is compatible with that of Doke’s [35]. However, it increases as the energy lowers. Previous measurements of this combined intrinsic resolution yielded very high values (6-8 ) [39], not compatible with our current measurement.
Furthermore, our work opens new possibilities for apparatus based on liquid xenon scintillators using scintillation only. In particular, it suggests that SSBs using LXe as a scintillator can be very competitive as building blocks in segmented calorimeters, with applications to nuclear and particle physics as well as PET technology. While there is ample room for technological development (eventually increasing the PDE of the SiPMs), our measurement clearly suggests that a LXe-SSB can compete and eventually outperform, in terms of energy resolution, those using other scintillators.

### v6

@@ I INTRODUCTION
All of the above crystals have either been used in, or actively pursued for, high energy and nuclear physics experiments. NaI was used in Crystal Ball, CsI in CLEO and BGO in L3. LYSO is one of the best candidates for future collider applications, due to its excellent properties (high light yield, fast decay time, high density) and radiation hardness [51]. LYSO is also the preferred scintillation material in modern Positron Emission Tomography (PET) scanners [48, 49, 17]. PET scanners benefit from a good energy resolution, because energy resolution allows one to reject a larger Compton scatter fraction, i.e., coincidences where one or both photons interact in the body of the patient and lose part of their energy. Those coincidences add noise to the image, since the direction of the gamma changes after the scattering.
@@
The common element in many nuclear and particle calorimeters and in modern PET scanners is a Segmented Scintillating Block, SSB (Fig. 1). An SSB is an array of contiguous cells filled with scintillating material. The light propagates inside each cell by reflection, until it reaches the readout sensor, normally a SiPM in modern devices, although PMTs and APDs were used in the past. The thickness of the cells depends on the application and varies between one and a few attenuation lengths. The SSB can be characterized by three parameters: energy resolution, which depends on the scintillator’s light yield, as well as the optical collection efficiency of the cell, the photodetection efficiency (PDE) of the SiPM, electronic noise, etc; time resolution, which is dominated by the scintillator’s decay time; and (transverse) spatial resolution, which can be tailored by choosing a suitable transverse size for the cell. As an example, in the 2 m long EXPLORER scanner [16, 9], the largest reported PET, the SSB material is LYSO, with cell area mm2 and thickness of 1.6 attenuation lengths, and the cells are read out by SiPMs. The number of individual cells exceeds . The energy resolution at 511 keV is 12% FWHM, the spatial resolution 3 mm and the time-of-flight (TOF) resolution is 500 ps.
@@
Liquid xenon (LXe) is also an excellent calorimeter [7]. Charged particles interact with LXe via both ionization and excitation. The excitation of xenon atoms results in strongly bound excited molecules (excimers) which, on de-excitation, emit VUV photons with a wavelength of 175 nm [24]. In addition, the xenon atoms are ionized. An external electric field can be used to carry away the ionization electrons, in which case two signals are available, one from scintillation, and one from ionization. In the absence of such electric field, most ionization electrons eventually recombine with positive xenon ions, and both channels (primary scintillation and recombination) contribute to the resulting scintillation yield of photons per MeV [13], which is a factor 1.5 larger than NaI and 1.7 larger than LYSO. Several measurements have found yield fractions of primary scintillation and recombination of roughly 0.26 and 0.74 for electrons of around 1 MeV of energy [32]. The primary scintillation is characterised by two decay constants ( ns, ns), while the decay time associated with recombination is ns [30]. Thus, LXe is, a priori, a brighter and (slightly) faster scintillator than LYSO, enabling improved energy resolution and time-of-flight performance, which in turn enhance the final image quality. However, its attenuation length is almost three times longer than that of LYSO and the fraction of photons which interact via the photoelectric effect (relevant for PET application) a factor of two less.
@@
LXe has been used extensively in particle and astroparticle physics. In particular, the MEG experiment [26] deployed a large calorimeter —800 litres volume filled with LXe—, with a thickness of 17 attenuation lengths and read out by 846 UV-sensitive photomultiplier tubes of 2-inch diameter. Other important examples are experiments searching for rare events, in particular neutrinoless double beta decay [6] and direct Dark Matter searches [7, 8]. Future experiments such as nEXO [2] and DARWIN [1] are also based on LXe. All these detectors are Time Projection Chambers, able to read the scintillation and ionization signals in LXe. Since the energy deposited by ionizing particles is distributed between both channels, the anti-correlation of the signals leads to a good energy resolution. For example, the XENON1T experiment measures 0.8% (all resolutions are expressed in terms of rms through the paper, unless stated otherwise) at 2.46 MeV [8] and the LZ experiment measures 0.67% at 2.61 MeV [40].
@@
The possibility of building a PET scanner based on LXe was proposed more than four decades ago [33] and explored by a number of subsequent works [14, 18, 21, 36]. The advantages of LXe compared to scintillating crystals are a large light yield, a fast decay time and the fact of being a liquid medium, which reduces the complexity of the system and increases its scalability. The disadvantage of having to add a cryogenic system to keep xenon in its liquid phase (the liquefaction temperature of xenon is -112∘C at 1 bar) is not problematic, since cryogenic systems reaching much lower temperatures are already in use in hospitals, for instance, to cool down magnets in magnetic resonance scanners. Two strategies were considered: a PET scanner based on scintillation only, and one that would combine both ionization and scintillation. Given the very high rate of PET scans, the second strategy may be unpractical, due to the long time (microseconds) needed to drift the charge. The LXe micro-PET prototype obtained an energy resolution of 4.3% [36] combining scintillation and charge, while the XEMIS1 prototype measured [25] with charge only. On the other hand, several groups tried prototypes based on pure scintillation (measured by PMTs), obtaining an energy resolution of 6.4–13.6% [21, 36, 15]. None of these results were compelling enough to consider LXe as a viable alternative to standard inorganic crystals, in particular LYSO. Recently, the possibility of building a LXe-PET using modern SiPM technology has been considered by our group. SiPMs offer improved energy resolution and time-of-flight performance, compactness enabling one-to-one crystal coupling for enhanced interaction reconstruction, and insensitivity to magnetic fields, allowing PET-MRI compatibility, being a significant improvement over earlier PMT-based PET scanners. [27, 28, 29, 42, 23, 43, 44].
@@
In the case of LXe, the intrinsic resolution comes up due to two components: the non-proportionality of the scintillation light yield with energy and the fluctuations in electron-ion recombination due to scape electrons. In 2002, Doke studied the non-proportionality of the scintillation light yield [22] resulting in in the region between 100 keV and 5 MeV (Fig. 2). His calculation is in good agreement with the modern Monte Carlo code describing the physics of LXe, NEST [47]. Specifically, at 511 keV, NEST predicts %.
Several experimental groups have reported energy resolution measurements in LXe for different gamma-rays, but none of these obtained the intrinsic resolution component related to the non-proportional light yield. For instance, a work from Yamashita et al. obtained an energy resolution of 12 at 122 keV and 7.1 at 662 keV [50], while Aprile et al. reported 5 at 122 keV [41]. The only attempt to go beyond simple energy resolution is that of Ni and collaborators, who estimated the intrinsic resolution as the combination of the non-proportionality of light yield and the electron-ion recombination fluctuations, obtaining 6-8 [39] for gamma rays between 122 keV and 662 keV. Although this value is not directly comparable to the result obtained by Doke, this result is almost a factor 5 worse.
In this paper, we present a new measurement of the energy resolution in LXe, using scintillation only. A value of the intrinsic resolution (non-proportionality + recombination fluctuations) is also shown. Our setup is conceptually identical to the SSB described above, and illustrated in Fig. 1, where the scintillating material is now LXe. The advantage of the SSB is its large light collection efficiency, as well as the availability of VUV-SiPMs with large PDE. This results in a large number of detected photoelectrons (pe), thus minimizing Poisson fluctuations and leading to a measurement close to .
@@ II SETUP
The active region of the setup consists of two LXe SSBs housed in a metal-sealed aluminum cube (F100-1ALU) from Vacom. The xenon is continuously recirculated in gas phase by a double diaphragm compressor and flows through a model PS4 MT15 R2 hot getter from Sigma Technologies. This process removes nitrogen, oxygen and water from the xenon, all of which quench xenon scintillation light [10]. The internal cube is housed within a large vacuum vessel so that the structure is thermally isolated while cooled. Said cooling is performed using a Sumitomo CH-110 cold head coupled to the cube via custom-made copper thermal links. Heat dissipation from the thermal links is minimized by wrapping them in polyethylene terephthalate foil. The cube and cooling machinery are visible in Fig. 4. The gas passes through a heat exchanger when entering and exiting the inner cube, which precools the gas before liquefaction for operation at 1.26 bar.
@@ III MONTE CARLO SIMULATION
A detailed simulation of our set-up has been developed using a Monte Carlo software application based on Geant4 [3]. The physical and optical properties of the materials in use are simulated thoroughly; in particular, the reflectivity of PTFE has been modeled as a Lambertian reflectivity with a value of and the value of the refractive index as a function of the wavelength has been extracted from Ref. [11]. The other physical parameters relevant for optical photon propagation are the Rayleigh scattering length (36 cm [46]) and the transparency and refractive index of the quartz window placed in front of the SiPMs to protect them (90 and 1.6, respectively). The photodetection efficiency provided by the sensor datasheet is also included as a function of the wavelength of the optical photons. We turn off any intrinsic fluctuations of the LXe response to gamma interactions, so that the amount of scintillation photons produced is proportional to the deposited energy (see more detailed information in Appendix A).
The simulation of a 22Na source placed in the calibration port provides a charge distribution with the photopeak centred in 5240 pe. Since the number of microcells in our SiPMs is 6162, our measurement is affected by saturation. In order to confirm that, we ran a second simulation, where each microcell of the SiPMs is simulated and the recovery time of our sensors (158 ns at -110∘C) taken into account, to model the loss of photons due to saturation; we found that in this case the photopeak is centered in 3752 pe, which implies a significant charge loss. The same simulations, with and without saturation effects, were carried out for two other available calibration sources, with lower energy gammas, namely 133Ba (provides gamma lines at 30.8, 81 and 356 keV) and 57Co (provides gamma line at 122 keV). In Fig. 6 the number of fired microcells versus the number of photons that should be detected in the case of negligible saturation is shown for MC events at all energies. The saturation curve provided by Hamamatsu [34] is superimposed on top of this distribution, showing the excellent match with the simulation. The positions of the photopeaks of the three calibration sources lie close to the Hamamatsu model.
@@ IV MEASUREMENT
The data acquisition output is a list of channels with integrated charge and corresponding timestamps. Once the channels are divided event by event, a first filter is applied to retain only coincidences, i.e., events where at least one sensor from each detection plane is present. This condition is applied only to data taken with the 22Na source. Given the cell structure of the SSBs, most of the charge is read out by the sensor coupled to the xenon volume where the interaction occurred, therefore, for all sources we single out the sensor with the larger detected charge for each plane.
Once the energy spectrum was obtained for the three sources, the ASIC non-perfectly linear response with energy due to the integrator was corrected and the energy spectrum in pe was obtained (see Fig. 7).
@@
where and are the best values for the Gaussian parameters coming from the fit. This procedure is applied to all the sensors in the SSB and their results are combined using a weighted mean, where each value is weighted by the inverse of its squared fit error (i.e., more precise measurements have a larger influence on the final resolution), thus yielding a final resolution of at 511 keV. The total uncertainty is calculated as the combination of two contributions: the statistical uncertainty from the fits and the spread across sensor resolutions.
The energy resolution is slightly better than the value predicted by MC without SiPM saturation, which is at 511 keV (see Fig. 9). The MC error includes the uncertainties in the optical modelling, such as the reflectivity of the teflon and the reflection and refraction between LXe and the quartz window that protects the SiPMs (see Fig. 3–bottom).
The low value of the energy resolution compared to MC is due to SiPM saturation, which has the effect of compressing the data and brings on an artificial better energy resolution. Therefore, correcting the data with the saturation curve provided by Hamamatsu, and fitting the peaks with the same procedure as before, an energy resolution of at 511 keV is obtained, showing compatibility with the MC result within . Fig. 10 shows the overlaid energy spectra of the experimental and MC data without saturation showing a good agreement. Fig. 11 presents the corrected energy resolutions for all the SiPMs in the reference SSB.
@@ V RESULTS
The energy resolution obtained in our work is shown in Fig.12, together with others’ measurements. It can be seen that our results are 2-3 times better than previous works using only scintillation. The only slightly better measurement is the one from Ref. [41], which only measured one peak.
Following Doke [22], the measured resolution, , can be seen as the quadratic sum of the the resolution related to the non-proportionality of scintillation light yield with energy, , the resolution due to Poisson variation in photon detection, and the resolution coming from the fluctuations of electron-ion recombination due to escape electrons, :
@@
Since our MC simulation does not include intrinsic fluctuations, we extract the intrinsic resolution (non-proportionality and fluctuations due to recombination), which we call , by subtracting the Poissonian width as determined by MC from the measured resolution.
@@ VI DISCUSSION
The intrinsic resolution that our procedure yields at 511 keV is compatible with that of Doke’s [22]. However, it increases as the energy lowers. Previous measurements of this intrinsic resolution yielded very high values (6-8 ) [39], not compatible with our current measurement.
Furthermore, our work opens new possibilities for apparatus based on liquid xenon scintillators using scintillation only. In particular, it suggests that SSBs using LXe as a scintillator can be very competitive as building blocks in segmented calorimeters, with applications to nuclear and particle physics as well as PET technology. While there is ample room for technological development eventually increasing the PDE of the SiPMs), our measurement clearly suggests that a LXe-SSB can compete and eventually outperform, in terms of energy resolution, those using other scintillators.
Appendix A Details of the Monte Carlo optical simulation
A dedicated Monte Carlo simulation of the experimental setup has been developed using the Geant4 toolkit [3]. This appendix provides additional details on the optical modeling assumptions and on the evaluation of the systematic uncertainties associated with the optical transport of scintillation photons in liquid xenon.
Optical surface model
The optical properties of the detector materials were implemented using the Geant4 unified model for optical photon interactions at surfaces. In this framework, the reflection properties of the PTFE were modeled assuming a purely diffuse (Lambertian) reflection component with a reflectivity of 98.
The choice of a Lambertian reflection model is motivated by several measurements reported in the literature. In particular, the study by Silva et al. [45] mentions more than 90 of the reflection is due to diffuse lobe at 0 degrees of incident angle, and more than 80 at 65 degrees.
Other studies have reported the presence of a specular component that increases for large angles of incidence (above –70°) [31]. However, the authors of that work also note a disagreement with previous results [4, 38] and attribute the discrepancy to a different sample surface preparation or surface degradation.
In our detector the PTFE surfaces are not optically polished or treated, therefore we assume the surface is rougher and the diffuse component may play a more important role. Furthermore, the implementation of an angle-dependent mixture of diffuse and specular reflection components is not straightforward within the standard Geant4 optical surface model. Considering both the experimental indications and the practical limitations of the implementation, we therefore adopted a purely diffuse reflection model as a reasonable approximation of the PTFE optical response.
Optical properties of the materials
The optical properties of liquid xenon relevant for photon transport were included in the simulation. The wavelength-dependent refractive index used in the simulation follows the measurements reported in Ref. [11]. The Rayleigh scattering length was set to 36 cm [46].
The absorption of scintillation photons in liquid xenon was neglected in the simulation. This approximation is justified by the very long attenuation length of liquid xenon at its own scintillation wavelength, which exceeds 10 m [10]. Since this value is much larger than the characteristic dimensions of our detector, the effect is expected to be negligible.
The optical interface between liquid xenon and the quartz window placed in front of the SiPMs was also modeled. The quartz window transparency was set to 90, and its refractive index was taken as 1.6, according to the manufacturer specifications. Reflection and refraction at the LXe–quartz interface were handled by the optical boundary processes implemented in Geant4. The wavelength-dependent photodetection efficiency of the SiPMs was implemented according to the sensor datasheet.
Evaluation of optical modeling uncertainties
Several optical parameters entering the simulation carry non-negligible uncertainties. To evaluate the impact of these uncertainties on the simulated detector response, we performed dedicated simulations in which the relevant parameters were varied within reasonable ranges.
In particular, the refractive index of the quartz window protecting the SiPMs was varied between 1.55 and 1.65 around the nominal value of 1.6. The reflectivity of the PTFE reflectors was varied between 98, as reported in the literature, and 99.99 to account for the high reflectivity of PTFE in the VUV range. In addition, simulations were performed both including and excluding the liquid xenon layer located behind the quartz window in order to evaluate the effect of this region on photon transport.
For each variation of these parameters the full simulation chain was repeated and the resulting changes in the predicted detector response were evaluated. The spread of the results obtained from these simulations is used to estimate the systematic uncertainty associated with the optical transport modeling and is included in the Monte Carlo results in section IV.